Stock S is expected to return 12% in a boom and 6% in a normal economy.

Stock T is expected to return 20% in a boom and 4% in a normal economy.

There is a probability of 40% that the economy will boom; otherwise, it will be normal.

What is the portfolio variance and standard deviation if 30% of the portfolio is invested in Stock S and 70% is invested in Stock T? Briefly discuss what this means to the stock.
Answer #1

Economy Portfolio Returns
Boom =30%*12%+70%*20%=17.60%

Normal =30%*6%+70%*4%=4.60%

Expected Returns=40%*17.60%+60%*4.60%=9.80%

Variance=0.40*(17.60%-9.80%)^2+0.60*(4.60%-9.80%)^2=0.004056

Standard Deviation=sqrt(0.004056)=6.37%